MA3000-D2-SU18A-Wiemann Han Quiz: LP04 Computational Quiz Time Remai This Question: 4 pts 11 of 11 (8 complete) Round to four decimal places as needed ) (Round to four decimal places as needed) Enter your answer in each of the answer boxes
Answer #1

\mathbf{SOLUTION}

\mathbf{given:}\lambda =5.6

\mathbf{formula:}P(X=x)=\frac{e^{-\lambda }*\lambda ^{x}}{x!}

\mathbf{{\color{Blue} a.}}\;P(X=1)=\frac{e^{-5.6}*5.6^{1}}{1!}={\color{Red} \mathbf{0.0207}}

\mathbf{{\color{Blue} b.}}\;P(X<1)\Rightarrow P(X=0)=\frac{e^{-5.6}*5.6^{0}}{0!}={\color{Red} \mathbf{0.0037}}

\mathbf{{\color{Blue} c.}}\;P(X>1)\Rightarrow1- P(X\leq 1)

P(X>1)=1-\left \{ P(X=0) +P(X=1)\right \}

P(X>1)=1-\left \{ 0.0037 +0.0207\right \}

P(X>1)={\color{Red} \mathbf{0.9756}}

\mathbf{{\color{Blue} d.}}\;P(X\leq 1)= P(X=0) +P(X=1)

P(X\leq 1)= 0.0037+0.0207={\color{Red} \mathbf{0.0244}}